OpenStudy (anonymous):

simplify cos^4x - sin^4x

Directrix (directrix):

cos^4x - sin^4x = (cos ^2 x + sin ^2 x) * (cos ^2 x - sin ^2 x). What does (cos ^2 x + sin ^2 x) equal? Hint: It is one of the identities.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what is it???

OpenStudy (aonz):

Do you know trid identities? How much do you know?

OpenStudy (anonymous):

im learning identities now. Dont know what trid means.. Its year 10 extension maths. I only know basic trig..

OpenStudy (aonz):

Sorry i meant trig... so what basic trig formulas do you know?

OpenStudy (anonymous):

sin cos tan

OpenStudy (aonz):

So you havent done sin2x and cos2x yet?

Directrix (directrix):

Look at the attached Pythagorean Identies: http://www.sosmath.com/trig/Trig5/trig5/trig5.html

Directrix (directrix):

cos^4x - sin^4x = (cos ^2 x + sin ^2 x) * (cos ^2 x - sin ^2 x) = ONE times (cos ^2 x - sin ^2 x)

OpenStudy (anonymous):

oh ok so i just need to remembe these identities... thx